OpenStudy (anonymous):

Which of the following represent the zeros of f(x) = 6x3 – 31x2 + 4x + 5 ?

OpenStudy (lalaly):

(x-5)(2x-1)(3x+1)=0

OpenStudy (anonymous):

f(5)=0 , so f(x)=(x-5)(6x^2-x-1), and the roots of (6x^2-x-1)=0 are x1=0.5 and x2=-1/3 Then f(x)=(x-5)(6x^2-x-1)=(x-5)6(x-0.5)(x+1/3) f(x)=(x-5)(2x-1)(3x+1)